OpenStudy (anonymous):

Solve. Check for extraneous solutions. Show work please. :) 5x = radical 30 - 5x

OpenStudy (anonymous):

25x^2=30-5x 5x^2+x-6=0 (x-1)(5x+6)=0 x=1,-6/5

OpenStudy (anonymous):

Thank you